### v1

@@ Appendix D Quantum reservoir characterization
Figure M5: Characterization of conditional displacements using an Anti-symmetric pulse from Ref. [1]. (a.) Pulse control schematic of characterization of conditional displacement. Here, we make use of number-splitting spectroscopy to characterize the state of the cavity after perform a conditional displacement on the state $|+\rangle|\mathrm{vacuum}\rangle$ conditioned on the qubit state. By post-selecting on the qubit state, we can evaluate the effectiveness of the conditional displacement. (b.) Number splitting spectroscopy conditioned on measuring the qubit in the excited state. A single parameter fit is used to capture the behavior of the state of the cavity as a function of amplitude. From the good agreement, we conclude that conditioned on measuring the qubit in the excited state, the cavity is displaced. (c.) Number splitting spectroscopy conditioned on measuring the qubit in the ground state. We see very limited change in the cavity state when measuring the qubit in the ground state.
Here, we provide the calibration of the CNOD unitary [1], one of the components of our reservoir unitary (Fig. M3). The CNOD protocol implements the following unitary
@@
Figure M5a displays the protocol for characterizing the anti-symmetric pulse. First, the qubit is unconditionally brought to the equator of the bloch sphere, with a wide-band $X_{\pi/2}$ pulse. After this, the anti-symmetric pulse acts on the cavity, followed by an qubit measurement, collapsing the cavity state to either $D(\alpha)|0\rangle$ or $|0\rangle$. After collapsing the state, we perform a number-splitting spectroscopy on the cavity. This is performed with a conditional $Y_{\pi}$, conditioned on the $k$th cavity Fock state [Gambetta2006, Schuster2007] followed by a second qubit measurement. By post-selecting on the first qubit measurement outcome, we can characterize the cavity state for each branch. Figure M5b and c show the number-splitting spectroscopy for the cavity state conditioned on the qubit being in the ground state vs excited, as a function of pulse amplitude. These curves are fitted with a single parameter scaling parameter that defines the relationship between pulse amplitude voltage and the amount of displacement $\alpha$.
@@
Figure M6: Characterization of input signal duration (a.) Diagram of control sequence to calibrate length of input signal duration toward implementing the geometric phase unitary in Eqs. 19-25. Here, a double conditional displacement is performed after sending the qubit to the equator of the Bloch sphere. After this, a variable delay is added before undoing that displacement. Finally, the qubit and the Fock distribution of the cavity is sampled using methods from Refs. [6, deng2023heisenberglimited]. (b.) Overlapping histograms showing Fock distribution of cavity state conditioned on qubit state as a function of delay. (c.) Cavity state overlap with the vacuum state as a function of delay. At a particular value of the delay, the two displacements interfere and cancel each other out.
@@
With our rotation gates and CNOD’s calibrated, we describe in this section the calibration of signal drives toward the implementation of Eqs. 19-25. We begin with a calibrating the duration of time our reservoir is exposed to the input signal. As discussed in Appendix C, calibrating this delay is crucial for a faithful implementation of the geometric phase detection unitary introduced in this work. While it may seem that this restriction in the signal duration is contrived and unrealistic in a real-world setting where the signal is unknown, we argue we can get around this by putting a commercial fast switch.
@@
Figure M6a schematically describes the experimental protocol for calculating the delay between the two CNOD pulses. Here, we effectively try to undo a double conditional displacement via second double conditional displacement. Due to the dispersive shift, after the first conditional displacement, the state of the cavity conditioned on the excited state of the qubit will start rotating with respect to the state of the cavity conditioned on the ground state. After a period of $2\pi/\chi$, the this will return to the same position as the start. Undoing the displacement at this point in time will send the cavity state to vacuum. Figure M6b shows the fock distribution of the cavity as a function of the waiting time, and Fig. M6c shows the cavity state overlap with the vacuum state as a function of the waiting time.
@@
Figure M7: Full geometric phase unitary calibration. (a.) Pulse control protocol for the unitary defined by Eqs 19-25. This protocol is identical to that of Fig. M3a. (b.) Overlapping histograms showing the probability over outcomes of measuring the qubit and oscillator-parity as a function of the phase of the input signal with $|\beta|=0.25$. As the phase is varied, the probability of the qubit being excited increases. (c.) The probability of measuring the qubit excited as the phase of the input displacement $\beta$ is varied, plotted for different values of $|\beta|$
@@
Next, we implement the full unitary given by Eqs. 19-25, where the section corresponding to the input data displacement (Eqs. 21 and 23) is given by the duration found in the results above. For this calibration, we implement the full unitary given by the diagram in Fig. M3a by varying the angle of the input displacement and looking at the dependence.
@@
Figure M7a shows the schematic overview of the calibration procedure. The geometric phase unitary is parameterized by a long displacement, whose angle we sweep. After the unitary we perform a qubit measurement, followed by a parity measurement. This calibration experiment is essentially identical to the time-independent reservoir computing experiments in terms of the control protocol. Here, instead of sending data from different distributions for the system to classify, we only vary the phase and amplitude of some input displacement so if we get the phase dependence we want.
@@
Figure M7b shows the distribution of measurement outcomes from measuring the qubit and the oscillator parity after the unitary is applied with $\alpha=1$ and $\beta=0.25$. As the angle of the input is swept, the qubit probability of the qubit being in found in the ground state shifts to being found in the excited state. This is more evident in Fig.M7c where we plot the probability of measuring the qubit in the excited state $P_{e}$ as a function of the phase of $\beta$ for different amplitudes of $\beta$. In comparison we find great qualitative agreement with the expected result $P_{e}=\cos(2|\alpha||\beta|\cos(\delta)+\pi/4)^{2}$, where $\delta=\arg(\alpha)-\arg(\beta)$ (see Eq.29), though we find an extra reduction in the dynamic range in $P_{e}$ for increasing $\beta$ due to qubit overheating.
@@
Figure M8: Qubit and oscillator-parity measurements in the quantum reservoir computer. For our reservoir construction, the state of the qubit is not generally known before a parity measurement. We apply feedback to change the parity measurement based on the preceding qubit measurement to faithfully capture the oscillator parity.
The qubit and parity measurements performed in this work are the standard pulse schemes used in many previous works, with one change. The typical procedure of measuring the parity of a cavity state is similar to a Ramsey experiment (and perhaps more closer still to a ‘qubit-revival’ experiment [chou2018teleported]), and importantly requires knowledge of the state of the qubit before the measurement is performed. In a quantum reservoir setting where measurement trajectories can be unknown, measuring the parity of the cavity is not straight-forward without post-selection or feedback. Here, since we perform a qubit measurement just before the parity measurement, we apply simple feedback that conditions the parity unitary on the measurement outcome of the preceding qubit measurement. The condition is such that the parity measurement outcome is now independent of the preceding measurement outcome. This reduces the order of correlations required to gain the same information: attaining the parity of the cavity only requires information about the parity measurement, whereas previously, second order correlations between the qubit and parity measurement was required. A further refinement to reduce trivial correlations in the measurement history would reset the qubit after the oscillator parity, however, due to limitations in the FPGA software, this was not implemented.
@@
Figure M9: Resonator-induced dephasing via pumping on the readout resonator (a.) Protocol for characterizing the effect of a readout pump on qubit decoherence: we perform a typically Ramsey experiment while populating the readout resonator to measure the $T_{2}$. (b.) Ramsey curves as a function of readout pump amplitude. These curves are fit using Eq. 38 to produce estimates of the qubit coherence. (c.) Extracted $T_{2}$ values for each of the curves in part (b)
Here we describe the experiment to reduce the qubit coherence time by pumping the readout resonator with photons during our reservoir experiments (see Fig. 2d). The calibration of this experiment involves performing standard a Ramsey $T_{2}$ experiment, modified with a pump on the readout resonator (Fig. M9a). Once populated, the resonator photons induce an dispersive shift, which, sends the qubit to the center of the Bloch sphere once the readout resonator is traced out. In principle, this interaction is coherent, and the qubit should see a revival; however, due to the leaky nature of the readout resonator by design, a coherent revival is not observed. As remarked in the end of Appendix B, this experiment required an auxilliary AWG line. Figure M1 denotes this as the ‘Readout Auxillary’ line.
Figure M9b shows the results of the Ramsey calibration with the readout pump on, for varying pump powers. We see a steady decrease in the qubit coherence time as the pump amplitude is increased as expected. The curves are fit to the equation
@@
where $\delta$ is an intentional detuning. Here, a Gaussian pulse was used as the readout pump. We expect that due to the construction of the reservoir, a flattop pulse may be more detrimental to the classification performance, since the Gaussian pulse has little amplitude during the CNOD unitaries shown in Fig. M3a. Finally, we note that the maximum $T_{2}$ shown in Fig. M9 differs from the value quoted in Table S1. After preliminary calibration data corresponding to those in Fig. M9, the experiments in Fig. 2c were performed, after which the qubit $T_{2}$ was suddenly lowered. However, all experiments presented in this manuscript, with the exception of Fig. M9, were performed where the qubit $T_{2}$ matched that of Table S1. Given the conclusion that the qubit $T_{2}$ did not impact classification accuracies until it approaches the time between measurements, we decided to include the higher quality data presented in Fig. M9, rather than the preliminary data used to calibrate the results in Fig. 2c.
Appendix E Machine learning with the quantum reservoir
@@
Figure M10: Comparison of feature vectors in spiral classification performance Here, classification accuracy on the spiral task is considered for different output encodings. Particularly, we compare included higher and higher correlations. $\vec{\mu}_{\leq p}$ describes a feature vector containing all central moments up to and including the $p$-th central moment (see text).
In this work, we use measurement correlations as the output feature vectors from which the trained linear layer of our reservoir performs the classification. In this section, we provide details in how these were constructed from measurement results, as well as motivations and comparisons with other output encodings. As described in the main text, measurements of our reservoir involve two measurements following every data input: a qubit measurement and a parity measurement. The qubit measurement, which follows just after the input unitary, either extracts information about the input displacement (if the signal is time-independent), or performs some nontrivial back-action on the oscillator state (see Fig. M4). The parity measurement, which follows the qubit measurement, will simply measure the parity of the cavity state post-qubit measurement, and collapse the oscillator state to either even or odd Fock states. It is worth pointing out that measurements of the parity are done with an entangling unitary starting with a known qubit state and then performing a regular qubit measurement (see Appendix D for details).
@@
where the number of indices of $\mu_{p}$ is equal to $p$. Here $x_{ni}$ is the $n$th repeated measurement result of observable $x$ labeled by $i$. In our setting, $i$ labels the $i$-th measurement in a sequence of correlated measurements before the system is reset. The expectation value $\langle x_{i}\rangle$ is taken over the shots $N_{\mathrm{shots}}$ – counting the number of system resets and repetitions. Faithful estimates of these moments are made with enough shots typically requiring on the order of a 1000 shots for the results presented in this manuscript.
@@
Figure M11: Second-order central moment (covariance) of quantum reservoir output over spiral dataset. These correlation matrices were generated from calculating the covariance over measurement outcomes in a reservoir run, then averaged over the entire dataset.
@@
The central moments of Eq. 39 are used in the construction of the output feature vector for the linear layer to perform the classification task. Specifically, the feature vector is generated by appending successively more and more central moments. We denote these appended feature vectors as $\vec{\mu}_{\leq p}$ for feature vectors containing up to $p$ central moments, e.g. $\vec{\mu}_{\leq 2}=[\vec{\mu}_{1},\mu_{2}]$ is a feature vector constructed from appending the covariance to the mean. The first order moment here is a vector to denote that the mean is taken over repetitions of different measurements, whereas the covariance is a matrix and thus is not denoted as a vector. Additionally, we only take the upper triangle of the covariance at most, since that contains all the independent degrees of freedom of the symmetric covariance matrix.
@@
Figure M12: Third-order central moment of quantum reservoir output over spiral dataset. These third-order central moments are plotted as an array of 2D matrices, such that the $i$-th column corresponds to the 2D matrix $\mu_{ijk}$. The two rows corresponds to the two different classes of signals. In the third order, one can begin to see differences between the two classes by eye.
@@
Figure M10 contains classification results on the spiral dataset (Fig. 2) as a function of the number of shots for the feature vectors $\vec{\mu}_{1},\vec{\mu}_{\leq 2}$ and $\vec{\mu}_{\leq 3}$. We see that our quantum reservoir has non-trivial third-order correlations and that the reservoir leverages these correlations to boost classification accuracy. The covariance matrix averaged over the entire spiral dataset is plotted in Fig. M11, and the third order correlations are plotted in Fig. M12 – plotted as a set of 2D matrices. In the third-order correlations in particular, we can begin to pick out by eye the differences in the two classes.
@@
In general, for arbitrary moments, the number of independent components is (${M+p-1}\choose{p}$, where $p$ is the order of the moment, and $M$ is the number of measurements. This construction generally allows us to construct feature vectors that are smaller than the sample probability over all possible measurement trajectories, which is $2^{M}$ dimensional. However, as can be seen in Fig. M11, there is yet redundant information even after taking only the symmetric part, specifically, that the information tends to be very local and that measurements far apart tend not to be correlated. This has the physical interpretation that while measurements are indeed correlated, even possessing higher-order correlations, this correlation tends to be local due to the finite memory of the system. This motivates us to further restrict our feature vector to only capture the essential local correlations.
@@
Figure M13 compares the classification performance of feature vectors generated with up to third-order moments, where we truncate the locality of the correlations. That is, the elements of the third order central moment $(\mu_{3})_{ijk}$ is set to zero if $|i-j|>d_{H}$ or $|i-k|>d_{H}$, for some integer $d_{H}$ we interpret as a Hamming distance. We note that including third-order correlations between measurements that are up to three ‘sites’ away nearly reproduces the classification accuracy of when you include all third-order central moments. Additionally, we compare the construction of feature vectors using truncated moments up to third-order with that of using the full sampled distribution as the feature vector. These last two statements were found to be true for all tasks presented in this paper. We used the truncated third-order correlations as the feature vector as the universal feature vector for all tasks presented.
@@
Figure M13: Spiral classification accuracy as a function of locality of up-to third order correlations Classification accuracy for the spiral task using up to third-order central moments truncating correlations to only include correlations up to $d_{H}$. We find that we can achieve high accuracy using local third-order correlations, saturating the accuracy when only keeping correlators up to $d_{H}\leq 3$. The performance of using the sample distribution is also compared which performs worse than using the central moments as output features, despite containing more information (see text).
@@
$y=W_{\text{train}}x+v_{\text{train}}$ (40)
where the largest of the $C$ elements of $y$ corresponded to the predicted class of the data point ($C$ is the number of classes). To train the linear layer, we chose between two different approaches: the pseudo-inverse method and back-propagation through a softmax function on the output.
First, we will describe the pseudo-inverse method. Let $X$ be a $N\times(R+1)$ matrix consisting of $R$-dimensional reservoir features generated for $N$ training points, with a row of 1’s appended (this is to compute both $W_{\text{train}}$ and $v_{\text{train}}$ at once). Let $Y$ be an $N\times C$ matrix consisting of $C-$dimensional column vectors that serve as labels for the training points such that $Y_{i,j}=1$ if $j$ corresponds to the class of the $i^{\text{th}}$ training point and zero otherwise. For an $\epsilon>0$, we construct $W^{{}^{\prime}}_{\text{train}}$ ($W_{\text{train}}$ appended with $v_{\text{train}}$) as :
@@
In our case, the value of $\epsilon$ was swept to maximize the accuracy of the classification. In the limit of $\epsilon\rightarrow 0$, pseudo-inverse matrix of Eq. 41 minimizes the mean squared error between $XW_{\text{train}}$ and $Y$, and so has been a popular choice for training the linear layer at the output of reservoirs [nakajima2018reservoir, LESNREVIEW2022, 2, scardapane2017]. However, our goal was to classify the input signals based on the largest element of the final output vector. Consequently, the linear layer that resulted in the lowest mean squared error with our labels was not always the linear layer that gave us the best accuracy.
For this reason, we also used a second training method for our linear layer. This approach used softmax and back-propagation using the automatic differentiation package from PyTorch [pytorch2019]. In this approach, the prediction vector $y$ from Eq. 40 is passed through the “Softmax” activation function:

### v2

@@ Appendix C Quantum reservoir characterization
Figure S5: Characterization of conditional displacements using an Anti-symmetric pulse from Ref. [1]. (a.) Pulse control schematic of characterization of conditional displacement. Here, we make use of number-splitting spectroscopy to characterize the state of the cavity after perform a conditional displacement on the state $|+\rangle|\mathrm{vacuum}\rangle$ conditioned on the qubit state. By post-selecting on the qubit state, we can evaluate the effectiveness of the conditional displacement. (b.) Number splitting spectroscopy conditioned on measuring the qubit in the excited state. A single parameter fit is used to capture the behavior of the state of the cavity as a function of amplitude. From the good agreement, we conclude that conditioned on measuring the qubit in the excited state, the cavity is displaced. (c.) Number splitting spectroscopy conditioned on measuring the qubit in the ground state. We see very limited change in the cavity state when measuring the qubit in the ground state.
Here, we provide the calibration of the CNOD unitary [1], one of the components of our reservoir unitary (Fig. S3). The CNOD protocol implements the following unitary
@@
Figure S5a displays the protocol for characterizing the anti-symmetric pulse. First, the qubit is unconditionally brought to the equator of the bloch sphere, with a wide-band $X_{\pi/2}$ pulse. After this, the anti-symmetric pulse acts on the cavity, followed by an qubit measurement, collapsing the cavity state to either $D(\alpha)|0\rangle$ or $|0\rangle$. After collapsing the state, we perform a number-splitting spectroscopy on the cavity. This is performed with a conditional $Y_{\pi}$, conditioned on the $k$th cavity Fock state [Gambetta2006, Schuster2007] followed by a second qubit measurement. By post-selecting on the first qubit measurement outcome, we can characterize the cavity state for each branch. Figure S5b and c show the number-splitting spectroscopy for the cavity state conditioned on the qubit being in the ground state vs excited, as a function of pulse amplitude. These curves are fitted with a single parameter scaling parameter that defines the relationship between pulse amplitude voltage and the amount of displacement $\alpha$.
@@
Figure S6: Characterization of input signal duration (a.) Diagram of control sequence to calibrate length of input signal duration toward implementing the geometric phase unitary in Eqs. 19-25. Here, a double conditional displacement is performed after sending the qubit to the equator of the Bloch sphere. After this, a variable delay is added before undoing that displacement. Finally, the qubit and the Fock distribution of the cavity is sampled using methods from Refs. [5, deng2023heisenberglimited]. (b.) Overlapping histograms showing Fock distribution of cavity state conditioned on qubit state as a function of delay. (c.) Cavity state overlap with the vacuum state as a function of delay. At a particular value of the delay, the two displacements interfere and cancel each other out.
@@
With our rotation gates and CNOD’s calibrated, we describe in this section the calibration of signal drives toward the implementation of Eqs. 19-25. We begin with a calibrating the duration of time our reservoir is exposed to the input signal. As discussed in Appendix B, calibrating this delay is crucial for a faithful implementation of the geometric phase detection unitary introduced in this work. While it may seem that this restriction in the input signal duration is contrived in a real-world setting where the signal is unknown, this restriction would be implemented via a fast switch that exposes our device to the unknown signal periodicially.
@@
Figure S6a schematically describes the experimental protocol for calculating the delay between the two CNOD pulses. Here, we effectively try to undo a double conditional displacement via second double conditional displacement. Due to the dispersive shift, after the first conditional displacement, the state of the cavity conditioned on the excited state of the qubit will start rotating with respect to the state of the cavity conditioned on the ground state. After a period of $2\pi/\chi$, the this will return to the same position as the start. Undoing the displacement at this point in time will send the cavity state to vacuum. Figure S6b shows the Fock distribution of the cavity as a function of the waiting time, and Fig. S6c shows the cavity state overlap with the vacuum state as a function of the waiting time.
@@
Figure S7: Full geometric phase unitary calibration. (a.) Pulse control protocol for the unitary defined by Eqs 19-25. This protocol is identical to that of Fig. S3a. (b.) Overlapping histograms showing the probability over outcomes of measuring the qubit and oscillator-parity as a function of the phase of the input signal with $|\beta|=0.25$. As the phase is varied, the probability of the qubit being excited increases. (c.) The probability of measuring the qubit excited as the phase of the input displacement $\beta$ is varied, plotted for different values of $|\beta|$
@@
Next, we implement the full unitary given by Eqs. 19-25, where the section corresponding to the input data displacement (Eqs. 21 and 23) is given by the duration found in the results above. For this calibration, we implement the full unitary given by the diagram in Fig. S3a by varying the angle of the input displacement and looking at the dependence.
@@
Figure S7a shows the schematic overview of the calibration procedure. The geometric phase unitary is parameterized by a long displacement, whose angle we sweep. After the unitary we perform a qubit measurement, followed by a parity measurement. This calibration experiment is essentially identical to the time-independent reservoir computing experiments in terms of the control protocol. Here, instead of sending data from different distributions for the system to classify, we only vary the phase and amplitude of some input displacement to get the phase dependence we want.
@@
Figure S7b shows the distribution of measurement outcomes from measuring the qubit and the oscillator parity after the unitary is applied with $\alpha=1$ and $\beta=0.25$. As the angle of the input is swept, the qubit probability of the qubit being in found in the ground state shifts to being found in the excited state. This is more evident in Fig.S7c where we plot the probability of measuring the qubit in the excited state $P_{e}$ as a function of the phase of $\beta$ for different amplitudes of $\beta$. In comparison we find great qualitative agreement with the expected result $P_{e}=\cos(2|\alpha||\beta|\cos(\delta)+\pi/4)^{2}$, where $\delta=\arg(\alpha)-\arg(\beta)$ (see Eq.29), though we find an extra reduction in the dynamic range in $P_{e}$ for increasing $\beta$ due to qubit overheating.
@@
Figure S8: Qubit and oscillator-parity measurements in the quantum reservoir computer. For our reservoir construction, the state of the qubit is not generally known before a parity measurement. We apply feedback to change the parity measurement based on the preceding qubit measurement to faithfully capture the oscillator parity.
The qubit and parity measurements performed in this work are the standard pulse schemes used in many previous works, with one change. The typical procedure of measuring the parity of a cavity state is similar to a Ramsey experiment (and perhaps more closer still to a ‘qubit-revival’ experiment [chou2018teleported]), and importantly requires knowledge of the state of the qubit before the measurement is performed. In a quantum reservoir setting where measurement trajectories can be unknown, measuring the parity of the cavity is not straight-forward without post-selection or feedback. Here, since we perform a qubit measurement just before the parity measurement, we apply simple feedback that conditions the parity unitary on the measurement outcome of the preceding qubit measurement. The condition is such that the parity measurement outcome is now independent of the preceding measurement outcome. This reduces the order of correlations required to gain the same information: attaining the parity of the cavity only requires information about the parity measurement, whereas previously, second-order correlations between the qubit and parity measurement was required. A further refinement to reduce trivial correlations in the measurement history would reset the qubit after the oscillator parity, however, due to limitations in the FPGA software, this was not implemented.
@@
Figure S9: Resonator-induced dephasing via pumping on the readout resonator (a.) Protocol for characterizing the effect of a readout pump on qubit decoherence: we perform a typically Ramsey experiment while populating the readout resonator to measure the $T_{2}$. (b.) Ramsey curves as a function of readout pump amplitude. These curves are fit using Eq. 38 to produce estimates of the qubit coherence. (c.) Extracted $T_{2}$ values for each of the curves in part (b)
Here we describe the experiment to reduce the qubit coherence time by pumping the readout resonator with photons during our reservoir experiments (see Fig. 2d). The calibration of this experiment involves performing a standard Ramsey $T_{2}$ experiment, modified with a pump on the readout resonator (Fig. S9a). Once populated, the resonator photons induce a dispersive shift, which sends the qubit to the center of the Bloch sphere once the readout resonator is traced out. In principle, this interaction is coherent, and the qubit should see a revival. However, due to the leaky nature of the readout resonator by design, a coherent revival is not observed. As remarked at the end of Appendix A, this experiment required an auxiliary AWG line. Figure S1 denotes this as the ‘Readout Auxiliary’ line.
Figure S9b shows the results of the Ramsey calibration with the readout pump on, for varying pump powers. We see a steady decrease in the qubit coherence time as the pump amplitude is increased as expected. The curves are fit to the equation
@@
where $\delta$ is an intentional detuning. Here, a Gaussian pulse was used as the readout pump. We expect that due to the construction of the reservoir, a flattop pulse may be more detrimental to the classification performance, since the Gaussian pulse has little amplitude during the CNOD unitaries shown in Fig. S3a. Finally, we note that the maximum $T_{2}$ shown in Fig. S9 differs from the value quoted in Table S1. After preliminary calibration data corresponding to those in Fig. S9, the experiments in Fig. 2c were performed, after which the qubit $T_{2}$ was suddenly lowered. However, all experiments presented in this manuscript, with the exception of Fig. S9, were performed where the qubit $T_{2}$ matched that of Table S1. Given the conclusion that the qubit $T_{2}$ does not impact classification accuracies until it approaches the time between measurements, we decided to include the higher quality data presented in Fig. S9, rather than the preliminary data used to calibrate the results in Fig. 2c.
Appendix D Machine learning with the quantum reservoir
@@
Figure S10: Comparison of feature vectors in spiral classification performance Here, classification accuracy on the spiral task is considered for different output encodings. Particularly, we compare included higher and higher correlations. $\vec{\mu}_{\leq p}$ describes a feature vector containing all central moments up to and including the $p$-th central moment (see text).
In this work, we use measurement correlations as the output feature vectors from which the trained linear layer of our reservoir performs the classification. In this section, we provide details in how these were constructed from measurement results, as well as motivations and comparisons with other output encodings. As described in the main text, measurements of our reservoir involve two measurements following every data input: a qubit measurement and a parity measurement. The qubit measurement, which follows just after the input unitary, either extracts information about the input displacement (if the signal is time-independent), or performs some nontrivial back-action on the oscillator state (see Fig. S4). The parity measurement, which follows the qubit measurement, will simply measure the parity of the cavity state post-qubit measurement, and collapse the oscillator state to either even or odd Fock states. It is worth pointing out that measurements of the parity are done with an entangling unitary starting with a known qubit state and then performing a regular qubit measurement (see Appendix C for details).
@@
where the number of indices of $\mu_{p}$ is equal to $p$. Here $x_{ni}$ is the $n$th repeated measurement result of observable $x_{i}$. In our setting, $i$ labels the $i$-th measurement in a sequence of correlated measurements before the system is reset. The expectation value $\langle x_{i}\rangle$ is taken over the shots $N_{\mathrm{shots}}$ – counting the number of system resets and repetitions. Faithful estimates of these moments typically require on the order of 1000 shots for the results presented in this manuscript.
@@
Figure S11: Second-order central moment (covariance) of quantum reservoir output over spiral dataset. These correlation matrices were generated from calculating the covariance over measurement outcomes in a reservoir run, then averaged over the entire dataset.
@@
The central moments of Eq. 39 are used in the construction of the output feature vector for the linear layer to perform the classification task. Specifically, the feature vector is generated by appending successively more and more central moments. We denote these appended feature vectors as $\vec{\mu}_{\leq p}$ for feature vectors containing up to $p$ central moments, e.g.
$\vec{\mu}_{\leq 2}$ $=[\vec{\mu}_{1},\mu_{2}]$
$=[\langle x_{0}\rangle,\langle x_{1}\rangle,\langle x_{2}\rangle,\ldots,\langle x_{0}x_{1}\rangle-\langle x_{0}\rangle\langle x_{1}\rangle,\langle x_{0}x_{2}\rangle-\langle x_{0}\rangle\langle x_{2}\rangle,\ldots,\langle x_{1}x_{2}\rangle-\langle x_{1}\rangle\langle x_{2}\rangle,\ldots]$
is a feature vector constructed from appending the flattened covariance to the mean. The first-order moment here is a vector to denote we take the mean over repetitions of different measurements, whereas the covariance is a matrix and thus is not denoted as a vector. Additionally, we only take the independent degrees of freedom of the symmetric covariance matrix equivalent to discarding one of the following redundant elements $\langle x_{i}x_{j}\rangle$ and $\langle x_{j}x_{i}\rangle$ for some integers $i,j$. In general, for arbitrary moments, the number of independent components for $M$ measurements is (${M+p-1}\choose{p}$, where $p$ is the order of the moment. For up to third-order central moments of $M=8$, this gives a total output feature dimension of $\mathrm{dim}(\vec{\mu}_{\leq 3})=8+36+120=164$. This output dimension for the results presented in the main text is further reduced as discussed in the following paragraphs below.
@@
Figure S12: Third-order central moment of quantum reservoir output over spiral dataset. These third-order central moments are plotted as an array of 2D matrices, such that the $i$-th column corresponds to the 2D matrix $\mu_{ijk}$. The two rows corresponds to the two different classes of signals. In the third order, one can begin to see differences between the two classes by eye.
@@
Figure S10 contains classification results on the spiral dataset (Fig. 2) as a function of the number of shots for the feature vectors $\vec{\mu}_{1},\vec{\mu}_{\leq 2}$ and $\vec{\mu}_{\leq 3}$. We see that our quantum reservoir has non-trivial third-order correlations and that the reservoir leverages these correlations to boost classification accuracy. The covariance matrix averaged over the entire spiral dataset is plotted in Fig. S11, and the third-order correlations are plotted in Fig. S12 – plotted as a set of 2D matrices. In the third-order correlations in particular, we can begin to pick out by eye the differences in the two classes.
@@
This construction generally allows us to construct feature vectors that are smaller than the probability distribution over all possible measurement trajectories, which is $2^{M}$ dimensional. However, as can be seen in Fig. S11, there is yet redundant information even after taking only the symmetric part - specifically, that the information tends to be very local and that measurements far apart tend not to be correlated. This has the physical interpretation that while measurements are indeed correlated, even possessing higher-order correlations, this correlation tends to be local due to the finite memory of the system. This motivates us to further restrict our feature vector to only capture the essential local correlations.
@@
Figure S13 compares the classification performance of feature vectors generated with up to third-order moments, where we truncate the locality of the correlations. That is, the elements of the third order central moment $(\mu_{3})_{ijk}$ are set to zero if $|i-j|>d_{H}$ or $|i-k|>d_{H}$, for some integer $d_{H}$. We note that including third-order correlations between measurements that are up to three ‘sites’ away nearly reproduces the classification accuracy of when you include all third-order central moments. Additionally, we compared the construction of feature vectors using truncated moments up to third-order with that of using the full sampled distribution as the feature vector and found that the former performed much better (Fig. S10). These last two statements were found to be true for all tasks presented in this paper. For $M=8$ measurements, the truncation of long range correlations further reduces the output feature size from 164 down to 94.
@@
Figure S13: Spiral classification accuracy as a function of locality of up-to third-order correlations Classification accuracy for the spiral task using up to third-order central moments truncating correlations to only include correlations up to $d_{H}$. We find that we can achieve high accuracy using local third-order correlations, saturating the accuracy when only keeping correlators up to $d_{H}\leq 3$. The performance of using the sample distribution is also compared which performs worse than using the central moments as output features, despite containing more information (see text).
@@
$y=x^{T}W_{\text{train}}+v_{\text{train}}$ (40)
where the largest of the $C$ elements of $y$ corresponded to the predicted class of the data point ($C$ is the number of classes). To train the linear layer, we chose between two different approaches: the pseudo-inverse method and back-propagation through a softmax function on the output. The two approaches optimize the linear layer over different loss landscapes. This is because our classification method is fundamentally discrete - i.e. we identify the class simply based on whichever output vector entry is the largest - so there is not a perfect correspondence between our loss and the classification inaccuracy.
First, we will describe the pseudo-inverse method. Let $X$ be a $N\times(R+1)$ matrix consisting of $R$-dimensional reservoir features generated for $N$ training points, with a column of 1’s appended (this is to compute both $W_{\text{train}}$ and $v_{\text{train}}$ at once). Let $Y$ be an $N\times C$ matrix consisting of $C-$dimensional row vectors that serve as labels for the training points such that $Y_{i,j}=1$ if $j$ corresponds to the class of the $i^{\text{th}}$ training point and zero otherwise. For an $\epsilon>0$, we construct $W^{{}^{\prime}}_{\text{train}}$ ($W_{\text{train}}$ appended with $v_{\text{train}}$) as :
@@
In our case, the value of $\epsilon$ was swept to maximize the accuracy of the classification. In the limit of $\epsilon\rightarrow 0$, the pseudo-inverse matrix of Eq. 41 is provably optimal for minimizing $||XW_{\text{train}}-Y||_{2}^{2}$ [leastsquarestextbook] up to numerical stability, and so has been a popular choice for training the linear layer at the output of reservoirs [nakajima2018reservoir, LESNREVIEW2022, 2, scardapane2017]. However, our goal was to classify the input signals based on the largest element of the final output vector. Consequently, the linear layer that resulted in the lowest mean squared error with our labels was not always the linear layer that gave us the best accuracy.
For this reason, we also used a second training method for our linear layer. This approach used softmax, a popular choice for classifiers in neural networks [banerjee2020] and back-propagation using the automatic differentiation package from PyTorch [pytorch2019]. Training through back-propagation with an optimizer is now necessary since an exact analytic solution to minimize the loss no longer exists, unlike the case of the pseudo-inverse. In this approach, the prediction vector $y$ from Eq. 40 is passed through the “Softmax” activation function: